# Probing supersolidity through excitations in a spin-orbit-coupled Bose-Einstein condensate

### v1

@@ IV.   Stripe compression excitations and location of the supersolid phase transition
Supersolids host a rich excitation spectrum originating from both their crystal-like and their superfluid character. Previous studies have revealed significant differences between platforms. Dipolar supersolids host crystal phonon-like modes [13, 15]. Instead, phonons are absent in cavity supersolids and multimode cavities are required to restore them [39]. This difference has led to controversial discussions on a supposed stiffness of the stripe pattern of spin-orbit-coupled supersolids, which were only recently clarified theoretically with the prediction of stripe compression modes akin to those of dipolar supersolids [26]. Here, we address this question experimentally. To this end, we prepare the system by a non-adiabatic ramp of the spin-orbit coupling strength $\Omega$, which is slow enough to remain in the lower dressed state, but sufficiently fast to excite the stripe pattern [29]. Figure 4(a) depicts the evolution of the modulation period $d$ over time for a system with $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$. Remarkably, we observe that $d$ is not constant, but instead displays a dynamical oscillation corresponding to a stripe compression mode. Its appearance proves that supersolids in spin-orbit-coupled BECs are not stiff. Observing this mode is a smoking gun of a fully dynamic supersolid, as it reveals the existence of a crystal phonon branch [26].
@@
In a momentum-space picture, this mode is caused by an out-of-phase oscillation of the $\ell$ and $r$ condensates in their dispersion minima, and within our effective mixture model thus corresponds to a spin-dipole mode. This mode disappears at the transition to the plane-wave phase. Its frequency has been theoretically predicted to reduce its value when approaching the supersolid stripe to plane-wave phase transition, reaching a minimum at the transition point $\Omega_{\mathrm{c}}$ [40, 41]. In the next series of experiments, we therefore focus on determining the critical value $\Omega_{\mathrm{c}}$ for a spin-orbit coupling detuning $\delta=0$ through the frequency of the compression mode. To measure its softening, it is essential to prepare the ground state of the system. In spin-orbit-coupled BECs, this is challenging due to the existence of very long-lived metastable states [20] with spin polarization $P\not=P_{0}$, where $P_{0}$ is the polarization of the ground state. In the limit of large system sizes our mixture model predicts that, for each value of $\Omega$, $P_{0}$ minimizes the stripe compression mode frequency $\omega_{\mathrm{SC}}$ [29]. We thus experimentally minimize $\omega_{\mathrm{SC}}$ as a function of polarization $P$ for different coupling strengths $\Omega$, which allows us to investigate the expected mode softening of the compression mode when approaching $\Omega_{\mathrm{c}}$. To enhance our detection in the regimes of weak coupling $\Omega$ and large spin polarization $P$, we exploit momentum space images to perform this measurement [29].
@@
Figure 4(b) summarizes our results. It displays the frequencies of the dipole and stripe compression modes as a function of the spin-orbit coupling strength $\Omega$, measured at the polarization $P$ that experimentally minimizes $\omega_{\mathrm{SC}}$, thus approximating the ground state polarization $P_{0}$. We observe that the frequency of both modes diminishes when increasing $\Omega$. For the dipole mode (red squares), it is due to the reduction of curvature of the dispersion relation at the minima (i.e., an increase of the effective mass [42, 43]). The frequency of the dipole mode vanishes at the plane-wave to single-minimum transition, which in the single-particle regime takes place at $\hbar\Omega=4\,E_{R}$ and has been investigated in previous works [35, 36]. It does not show particular features in the Raman coupling range explored here. In contrast, the frequency of the compression mode (blue circles), dramatically decreases when approaching the expected supersolid phase transition point $\Omega_{\mathrm{c}}$ (vertical black dashed line). The softening of the mode can also be clearly observed in the three exemplary oscillations included in the figure.
@@
We compare our experimental results with our theoretical prediction (solid lines) and find a good agreement. The theoretical lines have no fitting parameters and result from a method originally developed to analytically predict the mode frequencies in conventional mixtures [44], which we adapt to our mixture model [29]. In the vicinity of the phase transition point the experimental data deviates from the thermodynamic-limit theory due to finite size effects, which give a finite gap to the zero-energy modes. All measurements displayed in Fig. 4(b) are taken at a magnetic field $B=51.75\pm 0.02$ G, resulting in a supersolid phase transition located at $\hbar\Omega_{\mathrm{c}}=2.05\pm 0.08\,E_{\mathrm{R}}$ (vertical dashed line). By exploiting the rich Feshbach spectrum of 41K [32], we can furthermore tune the bare-state interactions and displace the phase transition point. The top panel of Fig. 4(c) summarizes the results of such measurements, which are in good agreement with the theoretical expectations (solid lines) and the predicted phase diagram of the system as a function of the magnetic field (bottom panel). This final measurement showcases the tunability of the supersolid phase transition in our system.
V.   Conclusion
@@
We have leveraged the tunable interaction properties of potassium atoms to realize a robust and high-contrast supersolid stripe phase in a spin-orbit-coupled BEC. We have imaged in situ its density modulation and have observed its rich dynamics. By observing a crystal compression mode, we have demonstrated that spin-orbit-coupled supersolid stripes are not stiff. We have used the frequency softening of this mode to locate the phase transition. Natural extensions of our work include the observation of other supersolid excitation modes, such as the crystal Higgs mode, and collective spin breathing [40, 33] and stripe angle oscillations [26], that could clarify how the spin degree of freedom further distinguishes spin-orbit-coupled supersolids from other platforms. Our system also opens the possibility of investigating the superfluid fraction across the phase transition [24]. Finally, by setting the effective interactions between the dressed condensates to attractive values ($a_{\ell r}<0$) it is possible to cancel the overall mean-field energy and reveal beyond mean-field effects. They are expected to stabilize quantum liquid droplets [45, 46] with supersolid properties [47, 48], thus providing a platform to systematically investigate the influence of quantum fluctuations on supersolidity.
@@ Acknowledgments
We acknowledge insightful discussions with S. Stringari, K. T. Geier, G. I. Martone, P. Hauke, W. Ketterle, I. B. Spielman, F. Ferlaino, and T. Donner. We also wish to thank T. Donner, W. Ketterle, P. Massignan, A. Rubio-Abadal and S. Stringari for a critical reading of the manuscript.
@@
We acknowledge funding from the European Union (ERC CoG-101003295 SuperComp), the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (LIGAS projects PID2020-112687GB-C21 at ICFO and PID2020112687GB-C22 at UAB, and Severo Ochoa CEX2019000910-S), the EU QuantERA project DYNAMITE (funded by MICN/AEI/ 10.13039/501100011033 and by the European Union NextGenerationEU/PRTR PCI2022-132919 (Grant No. 101017733)), Deutsche Forschungsgemeinschaft (Research Unit FOR2414, Project No. 277974659), Generalitat de Catalunya (AGAUR SGR 2021-SGR-01448 at ICFO and 2021-SGR-00138 at UAB, CERCA program), Fundació Cellex and Fundació Mir-Puig. C.S.C., S.H., and R.R. acknowledge support from the European Union (Marie Skłodowska-Curie–713729, Marie Skłodowska Curie–101149245 Epiquant, and Marie Skłodowska-Curie–101030630 UltraComp). V.B.M. acknowledges support from the Beatriu de Pinós Program and the Ministry of Research and Universities of the Government of Catalonia (Grant No. 2019-BP-00228). J.C. acknowledges support from the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and A.C. acknowledges support from the UAB Talent Research program during the early stages of this work.

### v2

@@ IV. Stripe compression excitations and location of the supersolid phase transition
Supersolids host a rich excitation spectrum originating from both their crystal-like and their superfluid character. Previous studies have revealed significant differences between platforms. Dipolar supersolids host crystal phonon-like modes [19, 21]. Instead, phonons are absent in cavity systems and multimode cavities are required to restore them [39]. This difference has led to controversial discussions on the stiffness of the stripe pattern of spin-orbit-coupled systems, which were only recently clarified theoretically with the prediction of stripe compression modes akin to those of dipolar supersolids [27]. Here, we address this question experimentally. To this end, we prepare the system by a non-adiabatic ramp of the spin-orbit coupling strength $\Omega$, which is slow enough to remain in the lower dressed state, but sufficiently fast to excite the stripe pattern [30]. Figure 4(a) depicts the evolution of the modulation period $d$ over time for a system with $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$. Notably, we observe that $d$ is not constant, but instead displays a dynamical oscillation corresponding to a stripe compression mode. This proves that the stripe phase of a spin-orbit-coupled BEC has a compressible crystal structure. Observing this mode is a signature of full-fledged supersolidity, as it reveals the existence of a crystal phonon branch [27].
@@
In a momentum-space picture, this mode is caused by an out-of-phase oscillation of the $\ell$ and $r$ condensates in their dispersion minima. Within our effective mixture model, it corresponds to a spin-dipole mode of the dressed states. Owing to its spin nature, it decouples from collective modes that involve instead the density degree of freedom, such as the breathing mode observed in Fig. 3(b) [30]. This mode disappears at the transition to the plane-wave phase, although it can be observed in out-of-equilibrium configurations [40]. Its frequency has been theoretically predicted to reduce when approaching the supersolid stripe to plane-wave phase transition, reaching a minimum at the transition point $\Omega_{\mathrm{c}}$ [41, 42]. In the next series of experiments, we therefore focus on determining the critical value $\Omega_{\mathrm{c}}$ for a spin-orbit coupling detuning $\delta=0$ through the frequency of the compression mode. To measure its softening, it is essential to prepare the ground state of the system. In spin-orbit-coupled BECs, this is challenging because of the existence of very long-lived metastable states [11] with spin polarization $P\not=P_{0}$, where $P_{0}$ is the polarization of the ground state. Such metastable states make it difficult to detect the phase transition point based only on the contrast of the stripes. Instead, in the limit of large system sizes our mixture model predicts that, for each value of $\Omega$, $P_{0}$ minimizes the stripe compression mode frequency $\omega_{\mathrm{SC}}$ [30]. We thus experimentally minimize $\omega_{\mathrm{SC}}$ as a function of polarization $P$ for different coupling strengths $\Omega$, which allows us to investigate the expected mode softening of the compression mode when approaching the critical point $\Omega_{\mathrm{c}}$ (Fig. S5) [30]. To enhance our detection in the regimes of weak coupling $\Omega$ or large spin polarization $P$, we exploit momentum space images to perform this measurement.
@@
Figure 4(b) summarizes our results. It displays the frequencies of the dipole and stripe compression modes as a function of the spin-orbit coupling strength $\Omega$, measured at the polarization $P$ that experimentally minimizes $\omega_{\mathrm{SC}}$, thus approximating the ground state polarization $P_{0}$. We observe that the frequency of both modes diminishes when increasing $\Omega$. For the dipole mode (red squares), this is caused by the reduction of curvature of the dispersion relation at the minima (i.e., an increase of the effective mass [43, 44]). The frequency of the dipole mode vanishes at the plane-wave to single-minimum transition, which in the single-particle regime takes place at $\hbar\Omega=4\,E_{R}$ and has been investigated in previous works [35, 36]. It does not show particular features in the Raman coupling range explored here. In contrast, the frequency of the compression mode (blue circles), dramatically decreases when approaching the expected supersolid phase transition point $\Omega_{\mathrm{c}}$ (vertical black dashed line). The softening of the mode can also be clearly observed in the three representative oscillations included in the figure.
@@
We compare our experimental results with our theoretical prediction (solid lines) and find a good agreement. The theoretical lines have no fitting parameters and result from a method originally developed to analytically predict the mode frequencies in conventional mixtures [45], which we adapt to our mixture model [30]. In the vicinity of the phase transition point the experimental data deviates from the thermodynamic-limit theory because of finite size effects, which give a finite gap to the zero-energy modes. All measurements displayed in Fig. 4(b) are taken at a magnetic field $B=51.75\pm 0.02$ G, resulting in a supersolid phase transition located at $\hbar\Omega_{\mathrm{c}}=2.05\pm 0.08\,E_{\mathrm{R}}$ (vertical dashed line). By exploiting the rich Feshbach spectrum of 41K [28], we can furthermore tune the bare-state interactions and displace the phase transition point. The top panel of Fig. 4(c) summarizes the results of such measurements, which are in good agreement with the theoretical expectations (solid lines) and the predicted phase diagram of the system as a function of the magnetic field (bottom panel). This final measurement showcases the tunability of the supersolid phase transition in our system.
V. Discussion and outlook
@@
Our realization of a high-contrast supersolid stripe phase in a spin–orbit-coupled BEC and the in situ observation of its crystal modes enables the observation of other excitations, in particular those involving the spin degree of freedom such as collective spin breathing [41, 33], stripe angle oscillations [27], and the Higgs amplitude mode of the translational symmetry breaking [18]. Although the existence of a crystal Higgs mode is a fundamental property of supersolids, in dipolar systems it remains unobserved because it couples to higher excited modes [46, 47]. In contrast, our mixture model suggests that in spin-orbit-coupled supersolids, the spin character of the Higgs mode keeps it decoupled from density excitations, making it more accessible. Besides their fundamental interest, investigating such excitations would clarify how the spin degree of freedom distinguishes spin-orbit-coupled supersolids from other platforms. Our system also opens the possibility of investigating the superfluid fraction across the supersolid phase transition [48], in analogy to very recent experiments with dipolar gases [24], but closer to the thermodynamic limit. Finally, by setting the effective interactions between the dressed condensates to attractive values ($a_{\ell r}<0$), it should be possible to cancel the overall mean-field energy and reveal beyond mean-field effects. These are expected to stabilize quantum liquid droplets [49, 50] with supersolid properties [51, 52], with lifetimes long enough to be observed. The ability to independently control the magnitude of beyond mean-field effects is a unique feature of spin-orbit-coupled supersolids, which should allow disentangling the effects of quantum fluctuations and supersolidity.
@@ Acknowledgments
We acknowledge insightful discussions with S. Stringari, K. T. Geier, G. I. Martone, P. Hauke, W. Ketterle, I. B. Spielman, F. Ferlaino, and T. Donner. We also wish to thank T. Donner, W. Ketterle, P. Massignan, A. Rubio-Abadal, and S. Stringari for a critical reading of the manuscript.
@@
We acknowledge funding from the European Union (ERC CoG-101003295 SuperComp), the Spanish Ministry of Science and Innovation MCIU/AEI/10.13039/501100011033 (projects LIGAS PID2020-112687GB-C21, MAPS PID2023-149988NB-C22, and Severo Ochoa CEX2024-001490-S at ICFO and projects LIGAS PID2020-112687GB-C22 and MAPS PID2023-149988NB-C21 PID2020112687GB-C22 at UAB), the EU QuantERA project DYNAMITE (funded by MICN/AEI/ 10.13039/501100011033 and by the European Union NextGenerationEU/PRTR PCI2022-132919 (Grant No. 101017733)), Deutsche Forschungsgemeinschaft (Research Unit FOR2414, Project No. 277974659), Generalitat de Catalunya (AGAUR SGR 2021-SGR-01448 at ICFO and 2021-SGR-00138 at UAB, CERCA program), Fundació Cellex and Fundació Mir-Puig. C.S.C., S.H., and R.R. acknowledge support from the European Union (Marie Skłodowska-Curie–713729, Marie Skłodowska Curie–101149245 Epiquant, and Marie Skłodowska-Curie–101030630 UltraComp). V.B.M. acknowledges support from the Beatriu de Pinós Program and the Ministry of Research and Universities of the Government of Catalonia (Grant No. 2019-BP-00228). J.C. acknowledges support from the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and A.C. acknowledges support from the UAB Talent Research program during the early stages of this work.
@@
Competing interests:
There are no competing interests to declare.
Data and materials availability:
The data shown in the main text and extended data figures are available from the CORA RDR repository [53].